OpenStudy (anonymous):

Can someone help me solve this? A study shows ceramic cups will increase by 2.5 percent next year, if 3,400,000 people use ceramic cups now how many people will use ceramic cups next year?

OpenStudy (johnweldon1993):

Well if 3400000 people use them now And that number will increase by 2.5 percent *or 0.025* then the # of people using them next year will be \[\large 3400000 + (3400000\times 0.025)\]

OpenStudy (anonymous):

I understand! Thank you for the help!

OpenStudy (anonymous):

@johnweldon1993 could I ask you one more question? If you given the points 1,3 and 2,3 which equation best fits the points? Y=x-2 Y=2x+1 Y=12x+4 Y=x+2

OpenStudy (johnweldon1993):

Sorry site is laggy lol So we have (1,3) and (2,3) ...if we plug in either pair to those equation choices...and see which one works for both pairs...we have the answer i.e for the first one y = x - 2 If we plug in (1,3) we have 3 = 1 - 2 *Not correct, so not that choice

OpenStudy (anonymous):

So how would I plug them in? For he second one would I do 2*1+1=3?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Would the answer be B?

OpenStudy (johnweldon1993):

Right...so our second choice is y = 2x + 1 So the points we have are (1,3) and (2,3) So checking (1,3) we have 3 = 2 + 1 ...good Now check (2,3) 3 = 2(2) + 1 ....not good...so nope not B

OpenStudy (anonymous):

I apologize I wrote the problem wrong, it was supposed to be 1,3, and 2,4

OpenStudy (anonymous):

I think D fits the answer best, because you do 1*x+2=3 and then 2*x+2=4